# Partonomy: Large Multimodal Models with Part-Level Visual Understanding

### v2

@@ 5.1 Explanatory Part Segmentation
In Table 2, PLUM outperforms LISA and GLaMM on all three part-segmentation question types in the zero-shot setting. We attribute this gain to (i) span-level constraints that keep pre-trained textual semantics intact and (ii) our mask-feedback loop, which refines each mask using its visual history. By contrast, Table 3 shows only marginal gaps in text-only metrics (P, R, Acc.). Because each answer choice shares extensive lexical overlap, the probability distribution across choices is nearly flat, dampening model-to-model variance on this challenging task.
@@
LISA-13B 5.9 7.0 7.1 7.5 6.1 7.1 5.7 6.6 6.0 6.8
PixelLM-13B 6.8 8.4 4.4 4.8 4.2 4.8 4.6 5.4 6.3 7.8
@@
GLaMM (ft) (rasheed2024glamm, ) 36.6 38.8 40.3 42.1 33.6 34.8 36.1 38.5 35.7 38.0
PLUM (ft) 36.2 41.6 42.1 45.9 33.0 39.4 36.7 40.8 36.2 39.8
@@
The Part-Whole Reasoning mask prediction results shown in Table 2 show that predicting the object label first prior to generating the part segmentation masks leads to better mask prediction performance (Whole-to-Part). This result is consistent across different segmenting LMMs, suggesting that part mask prediction benefits from object label conditioning. Similarly, the increase in object accuracy seen by predicting parts prior to the object label (Part-to-Whole) underscores the utility in jointly predicting object labels and parts.
@@ 5.2 Additional Downstream Tasks and Ablation Study
OVSeg liang2023open 28.5 LISA-13B 48.9
X-Decoder zou2023generalized 22.6 LISA-13B (ft) 56.2
SEEM zou2023segment 25.5 LISA-7B lai2024lisa 44.4
Grounded-SAM ren2024grounded 26.0 LISA-7B (ft) 52.9
@@
LISA-13B (lai2024lisa, ) 1.58 ($-93.1$%) 1.33 ($-97.4$%) 1.87 ($-94.1$%)
PixelLM-13B (ren2024pixellm, ) 8.37 ($-63.4$%) 21.72 ($-57.6$%) 15.29 ($-51.9$%)
@@
To investigate whether Plum’s removal of special tokens mitigates distribution shift from LLaVA’s pre-trained representations, we compare Plum’s performance on VQA (singh2019towards, ; hudson2019gqa, ; li2023evaluating, ) tasks to those of models which use [SEG] tokens and to that of the base LLaVA model Table 5. To our surprise, the performance of state-of-the-art segmenting LMMs deteriorates significantly on VQA and visual hallucination li2023evaluating tasks. Through inspection, we find that LISA tends to only generates irrelevant [SEG] tokens. While PixelLM performs somewhat better, it still suffers due to its use of multiple specialized tokens (ren2024pixellm, ). In contrast, Plum outperforms the other segmenting LMMs, even outperforming the LLaVA backbone on 2/3 tasks. This strong performance demonstrates that our paradigm of segmentation through span extraction preserves the visual reasoning capacity of the LMM.
@@
Removing the iterative mask-feedback loop lowers micro-gIoU by -6.5% and macro-gIoU by -6.3%, though the span-based tag extractor alone still beats the LISA-13B baseline by +6.7% on macro-gIoU. When both components are active, Plum tops LISA by +2.7% (micro) and +13.0% (macro), showing that tag extraction broadens long-tail coverage while feedback refines segmentation accuracy.
@@ 6 Limitations and Broader Impacts
Our work addresses an important yet underexplored problem in fine-grained, part-level understanding. While our proposed Partonomy-Core has the largest number of part labels out of part-grounding benchmarks to date—covering 862 part categories across 534 objects—it still omits some rare, domain-specific concepts and does not have the variety in object labels that he2022partimagenet has. Extending the dataset to cover a richer set of concepts (e.g. by incorporating PartImageNet++ into the Explanatory Part Segmentation task) would strengthen part-level understanding of LMMs. Although Plum overcomes key architectural challenges of existing segmenting LMMs through BIO tagging and a FiLM-based feedback loop, segmenting small or ambiguous parts remain challenging, and the model may not scale seamlessly to high-resolution images (a challenge shared by all LMMs using CLIP). Future work could explore multi-scale feature integration to close these gaps.
@@
Fine-grained part grounding is a cornerstone for compositional visual reasoning and has immediate implications for real-world systems—such as fine-grained robotic manipulation, assistive robotics, and automated inspection—where misunderstanding an object’s parts can lead to costly errors or safety hazards. By introducing Partonomy and Plum, we aim to catalyze progress towards more reliable and explainable LMMs. At the same time, our dataset may reflect biases present in its source images and annotations (e.g., cultural contexts), so rigorous bias auditing and inclusive data collection should accompany future expansions. Finally, the computational cost of training large segmenting LMMs underscores the need for research into more efficient architectures and sustainable training practices. We hope our benchmark, analysis of architectural shortcomings, and Plum design will inspire the community to pursue robust, efficient, and responsible part-grounding methods that generalize across domains and safeguard real-world deployments.
@@ A.1 Model Setup
Hyperparameter PLUM LISA PixelLM GLaMM Backbone Language model LLaMA-2-13B LLaMA-2-13B LLaMA-2-13B LLaMA-2-13B Vision tower CLIP ViT-L/14 CLIP ViT-L/14 CLIP ViT-L/14 CLIP ViT-L/14 Mask decoder SAM ViT-H SAM ViT-H Conv-U-Net SAM ViT-H I/O and training schedule Input resolution ($\mathrm{px}^{2}$) $1024^{2}$ $1024^{2}$ $1024^{2}$ $1024^{2}$ Max text length 512 512 512 512 Precision bf16 bf16 bf16 bf16 Epochs 25 (0-shot) | 4 (ft) / | 4 (ft) / | 4 (ft) / | 4 (ft) Batch size 6 6 6 6 Grad. accumulation 10 10 10 10 Effective batch 10×10$\times$10 × bsz $\times$ GPU 10×10$\times$10 × bsz $\times$ GPU 10×10$\times$10 × bsz $\times$ GPU 10×10$\times$10 × bsz $\times$ GPU Optimizer Optimizer AdamW AdamW AdamW AdamW Learning rate $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$ Betas (0.9, 0.95) id. id. id. Weight decay 0 0 0 0 Gradient clip 1.0 1.0 1.0 1.0 Loss weights $\lambda_{\mathrm{CE}}$ (LM) 1.0 1.0 1.0 1.0 $\lambda_{\mathrm{seg}}$ (Dice/FTL) 8.0  (FTL) 0.5 (Dice) 0.5 (Dice) 0.5 (Dice) $\lambda_{\mathrm{BCE}}$ (mask) 2.0 2.0 2.0 2.0 $\lambda_{\mathrm{KL}}$ 0.1 $\lambda_{\mathrm{cls}}$ (BIO) 2.0 Dice scale factor $10^{3}$ $10^{3}$ $10^{3}$ $10^{3}$ FTL $(\alpha,\beta)$ (0.7, 0.3) Additional Modules BIO span tagger Bidirectional encoder 2048 Feedback Loop (Temporal Mask Pooler)) Trainable SAM parts decoder $+$ prompt-enc. decoder decoder LoRA on LM (q,v) $r{=}8$, $\alpha{=}16$, $p{=}0.05$ id. id. id.
@@
Training uses DeepSpeed ZeRO-2 with bf16 precision, a per-GPU batch of 6, and gradient_accumulation_steps$=10$ (effective batch $10\times\text{bsz}\times\!N_{\text{GPU}}$). Weights are updated by AdamW ($\beta{=}(0.9,0.95)$, no weight-decay) with a peak learning-rate of $3\times 10^{-4}$, linearly warmed up for the first $100$ optimization steps and clipped to a global norm of $1.0$ thereafter.
We freeze the CLIP vision tower and the MM projection layer; all other components are trainable. The LLaMA decoder receives LoRA adapters on q_proj,v_proj ($r{=}8$, $\alpha{=}16$, dropout $0.05$). On the vision side we finetune the SAM mask-decoder and, when specified, the prompt-encoder. PLUM’s additional heads, the bidirectional BIO encoder and the Temporal Mask Pooler for mask feedback loop, are always optimized.
@@ Appendix B Partonomy Construction
For Partonomy-Core, we first construct the object-part hierarchy as illustrated in Figure 6. The hierarchy starts with a set of root-level, superordinate object categories (e.g., airplane, kitchen tools), where each category contains a set of intermediate object-level concepts (e.g., agricultural airplane, coffee maker) that fall under each superordinate category, and part-level concepts which compose the leaf nodes of each object-level concept. We manually define 10 distinct superordinate concept categories, ranging from generic concepts, e.g., vehicle, office supplies, to more complex concepts such as geography [^3] We thank MIT Lincoln Lab and DARPA for developing an initial category and concept ontology as part of the DARPA ECOLE effort, which we refine and expand to construct our part-centric ontology.. There are a total of 534 object-level concepts (e.g., banana boat under the boats category) These concepts contain 1976 concept-specific parts (e.g. biplane-wing, or 862 unique part types, each appearing in 1.6 object categories on average. Table 7 shows the full list of object categories with example object labels.
@@
Dataset statistics for our merged Partonomy dataset, containing all subsets (Pascal Part, PartImageNet, PACO, and Partonomy-Core) can be found in Appendix B.
@@
{tblr}
colspec = p2.9cm c c c c c, row1 = font=, width = Datasets # Object Labels # Part Labels # Object-Part Labels # Images # Seg. Masks
@@ C.2 Effect of fine-tuning on Partonomy
Fine-tuning on only the Partonomy training split yields double-digit improvements across all models, yet three consistent patterns emerge. PLUM (ft) attains the strongest macro-gIoU scores. It tops all three datasets in Identification and Intersection (e.g., 80.3 and 77.4 on PartImageNet), and remains within $\leq$ 2.0 micro-gIoU of the best performer, demonstrating excellent part recall on rare classes. GLaMM (ft) is highly competitive on micro metrics. GLaMM (ft) slightly edges out PLUM on Difference-micro for PartImageNet and PascalParts, likely because its DETR-style mask decoder better captures dominant foreground regions. Nevertheless, the gap in macro-gIoU ($\geq$ 4.0 gIoU) shows that GLaMM still under-segments uncommon parts. PixelLM-13B (ft) saturates early without intersection gains. Although PixelLM enjoys the largest absolute jump on PACO_LVIS (39.0 micro-gIoU in Identification), its Intersection scores plateau, implying limited benefit when parts touch or overlap.

### v3

@@ 5.1 Explanatory Part Segmentation
In Table 2, PLUM outperforms LISA and GLaMM on all three part-segmentation question types in the zero-shot setting. We attribute this gain to (i) span-level constraints that keep pre-trained textual semantics intact and (ii) our mask-feedback loop, which refines each mask using its visual history. By contrast, Table 3 shows only marginal gaps in text-only metrics (P, R, Acc.). Partonomy’s answer choices intentionally contain extensive lexical overlap, demonstrating the language models’ difficulties in comparing similar answer choices.
@@
Grounded SAM 2 (gt) (ren2024grounded, ) 13.6 16.8 20.6 23.6 14.3 17.1
LLaVAv1.5 + HIPIE (liu2024llavav1.5, ; wang2023hipie, )
SegLLM 29.6 32.4 32.2 33.8 28.5 30.7 29.4 32.3 29.3 32.2
LLaVA + SAM-style Mask Decoder (liu2023llava, ; kirillov2023segment, )
LISA-13B (lai2024lisa, ) 5.9 7.0 7.1 7.5 6.1 7.1 5.7 6.6 6.0 6.8
PixelLM-13B (ren2024pixellm, ) 6.8 8.4 4.4 4.8 4.2 4.8 4.6 5.4 6.3 7.8
@@
GLaMM (ft) 36.6 38.8 40.3 42.1 33.6 34.8 36.1 38.5 35.7 38.0
PLUM-13B (ft) 36.2 41.6 42.1 45.9 33.0 39.4 36.7 40.8 36.2 39.8
@@
The Part-Whole Reasoning results of Table˜2 show that knowing the object label prior to part segmentation leads to better mask prediction performance—the pretrained models obtain higher Whole2Part than Part2Whole scores. This suggests that part mask prediction benefits from object label conditioning. The advantage obtained by object conditioning evaporates once the models have been trained on sufficient part data, however, as shown by the finetuned models. Similarly, in Table˜3 the models’ increase in object accuracy after conditioning on the object’s parts, and their increase in part accuracy after conditioning on the object label, underscores the utility of jointly predicting object labels and parts.
@@ 5.2 Additional Downstream Tasks and Ablation Study
OVSeg liang2023open 28.5 LISA-7B lai2024lisa 44.4
X-Decoder zou2023generalized 22.6 LISA-7B (ft) 52.9
SEEM zou2023segment 25.5 LISA-13B 48.9
Grounded-SAM ren2024grounded 26.0 LISA-13B (ft) 56.2
@@
LISA-13B (lai2024lisa, ) 1.58 (93.1-93.1%) 1.33 (97.4-97.4%) 1.87 (94.1-94.1%)
PixelLM-13B (ren2024pixellm, ) 8.37 (63.4-63.4%) 21.72 (57.6-57.6%) 15.29 (51.9-51.9%)
@@
To investigate whether Plum’s removal of special tokens mitigates distribution shift from LLaVA’s pre-trained representations, we compare Plum’s performance on VQA (singh2019towards, ; hudson2019gqa, ; li2023evaluating, ) tasks to those of models which use [SEG] tokens and to that of the base LLaVA model Table˜5. To our surprise, the performance of state-of-the-art segmenting LMMs deteriorates significantly on VQA and visual hallucination li2023evaluating tasks. Through inspection, we find that LISA tends to only generates irrelevant [SEG] tokens. While PixelLM performs somewhat better, it still suffers due to its use of multiple specialized tokens (ren2024pixellm, ). In contrast, Plum outperforms the other segmenting LMMs, even outperforming the LLaVA backbone on 2/3 tasks. This strong performance demonstrates that our paradigm of segmentation through span extraction preserves the visual reasoning capacity of the LMM.
@@
Removing the iterative mask-feedback loop lowers micro-gIoU by 9.6% and macro-gIoU by 8%, though the span-based tag extractor alone still beats the LISA-13B baseline by 8.4% macro-gIoU. When both components are active, Plum tops LISA by 3.5% (micro) and 20% (macro), showing that tag extraction broadens long-tail coverage while feedback refines segmentation accuracy.
@@ 6 Limitations and Broader Impacts
Our work advances fine-grained, part-level understanding but still faces several constraints. Although Partonomy-Core includes the largest number of part labels to date, with 862 categories across 534 objects, it omits some rare or domain-specific parts and lacks the object diversity of he2022partimagenet . Expanding its coverage would further enhance LMM comprehension. While Plum mitigates major architectural limitations of prior segmenting LMMs via BIO tagging and a FiLM-based feedback loop, it still struggles with small or ambiguous parts and may not scale efficiently to high-resolution imagery.
@@
Part grounding is crucial for compositional visual reasoning and safety-critical domains such as robotic manipulation, assistive systems, and automated inspection. Such scenarios require accurate and interpretable part understanding, with mistakes in identifying and the grounding of parts leading to catastrophic consequences. By introducing Partonomy and Plum, we aim to foster more reliable and interpretable LMMs. The high computational cost of large segmenting LMMs also underscores the need for more efficient, sustainable architectures. We hope our benchmark and analyses inspire continued research toward robust, efficient, and responsible part-grounding methods.
@@ 8 Acknowledgements
The authors thank Steven Gomez and MIT Lincoln Laboratory for proposing an initial set of part-centric concepts as part of the DARPA ECOLE Program. We also thank the numerous annotators who contributed to the Partonomy dataset.
@@ A.1 Model Setup
Hyperparameter PLUM LISA PixelLM GLaMM Backbone Language model LLaMA-2-13B LLaMA-2-13B LLaMA-2-13B LLaMA-2-13B Vision tower CLIP ViT-L/14 CLIP ViT-L/14 CLIP ViT-L/14 CLIP ViT-L/14 Mask decoder SAM ViT-H SAM ViT-H Conv-U-Net SAM ViT-H I/O and training schedule Input resolution ($\mathrm{px}^{2}$) $1024^{2}$ $1024^{2}$ $1024^{2}$ $1024^{2}$ Max text length 512 512 512 512 Precision bf16 bf16 bf16 bf16 Epochs 25 (0-shot) | 4 (ft) / | 4 (ft) / | 4 (ft) / | 4 (ft) Batch size 6 6 6 6 Grad. accumulation 10 10 10 10 Effective batch $10\times$ bsz $\times$ GPU $10\times$ bsz $\times$ GPU $10\times$ bsz $\times$ GPU $10\times$ bsz $\times$ GPU Optimizer Optimizer AdamW AdamW AdamW AdamW Learning rate $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$ $3{\times}10^{-4}$ Betas (0.9, 0.95) id. id. id. Weight decay 0 0 0 0 Gradient clip 1.0 1.0 1.0 1.0 Loss weights $\lambda_{\mathrm{CE}}$ (LM) 1.0 1.0 1.0 1.0 $\lambda_{\mathrm{seg}}$ (Dice/FTL) 8.0  (FTL) 0.5 (Dice) 0.5 (Dice) 0.5 (Dice) $\lambda_{\mathrm{BCE}}$ (mask) 2.0 2.0 2.0 2.0 $\lambda_{\mathrm{KL}}$ 0.1 $\lambda_{\mathrm{cls}}$ (BIO) 2.0 Dice scale factor $10^{3}$ $10^{3}$ $10^{3}$ $10^{3}$ FTL $(\alpha,\beta)$ (0.7, 0.3) Additional Modules BIO span tagger Bidirectional encoder 2048 Feedback Loop (Temporal Mask Pooler)) Trainable SAM parts decoder ++ prompt-enc. decoder decoder LoRA on LM (q,v) r=8r{=}8, $\alpha{=}16$, p=0.05p{=}0.05 id. id. id.
@@
Training uses DeepSpeed ZeRO-2 with bf16 precision, a per-GPU batch of 6, and gradient_accumulation_steps=10=10 (effective batch $10\times\text{bsz}\times\!N_{\text{GPU}}$). Weights are updated by AdamW ($\beta{=}(0.9,0.95)$, no weight-decay) with a peak learning-rate of $3\times 10^{-4}$, linearly warmed up for the first $100$ optimization steps and clipped to a global norm of $1.0$ thereafter.
We freeze the CLIP vision tower and the MM projection layer; all other components are trainable. The LLaMA decoder receives LoRA adapters on q_proj,v_proj (r=8r{=}8, $\alpha{=}16$, dropout $0.05$). On the vision side we finetune the SAM mask-decoder and, when specified, the prompt-encoder. PLUM’s additional heads, the bidirectional BIO encoder and the Temporal Mask Pooler for mask feedback loop, are always optimized.
@@ Appendix B Partonomy Construction
For Partonomy-Core, we first construct the object-part hierarchy as illustrated in Figure 6. The hierarchy starts with a set of root-level, superordinate object categories (e.g., airplane, kitchen tools), where each category contains a set of intermediate object-level concepts (e.g., agricultural airplane, coffee maker) that fall under each superordinate category, and part-level concepts which compose the leaf nodes of each object-level concept. We manually define 10 distinct superordinate concept categories, ranging from generic concepts, e.g., vehicle, office supplies, to more complex concepts such as geography [^5] We thank MIT Lincoln Lab and DARPA for developing an initial category and concept ontology as part of the DARPA ECOLE effort, which we refine and expand to construct our part-centric ontology.. There are a total of 534 object-level concepts (e.g., banana boat under the boats category) These concepts contain 1976 concept-specific parts (e.g. biplane-wing, or 862 unique part types, each appearing in 1.6 object categories on average. Table˜7 shows the full list of object categories with example object labels.
@@
Dataset statistics for our merged Partonomy dataset, containing all subsets (Pascal Part, PartImageNet, PACO, and Partonomy-Core) can be found in Table˜8.
@@
Datasets # Object Labels # Part Labels # Object-Part Labels # Images # Seg. Masks
@@ C.2 Effect of fine-tuning on Partonomy
Fine-tuning on only the Partonomy training split yields double-digit improvements across our baselines, with three consistent patterns emerging. PLUM (ft) attains the strongest macro-gIoU scores. It tops all three datasets in Identification and Intersection (e.g., 80.3 and 77.4 on PartImageNet), and remains within $\leq$ 2.0 micro-gIoU of the best performer, demonstrating excellent part recall on rare classes. GLaMM (ft) is highly competitive on micro metrics. GLaMM (ft) slightly edges out PLUM on Difference-micro for PartImageNet and PascalParts. Nevertheless, the gap in macro-gIoU ($\geq$ 4.0 gIoU) shows that GLaMM still under-segments uncommon parts. PixelLM-13B (ft) saturates early without Intersection gains.
@@ C.3 Dataset difficulty and domain shift
C.4 Additional Ablation Studies
In addition to the main experiments, we conducted several ablations and follow-up analyses. These studies validate PLUM’s architectural choices, loss formulation, robustness, and dataset design.
Bidirectional attention for BIO tagging.
We confirmed that bidirectional attention is critical for accurate span extraction. Removing it severely degraded “I” tag accuracy on both Partonomy–PartImageNet and RefCOCO.
Table 12: Bidirectional attention ablation. Removing bidirectionality collapses span tagging performance.
Model B-Acc I-Acc O-Acc
PLUM (bi) 98.59 87.32 99.98
PLUM (no bi) 100.00 15.86 99.78
PLUM (bi, RefCOCO) 99.98 99.87 100.00
PLUM (no bi, RefCOCO) 6.68 4.92 99.98
Loss function comparison.
To isolate the impact of the Focal–Tversky loss (FTL), we retrained PLUM with Dice loss under identical conditions.
Table 13: DICE vs. Focal–Tversky. PLUM’s gains stem from its architecture choices; FTL provides marginal improvement over DICE.
Loss micro-gIoU macro-gIoU B-Acc I-Acc O-Acc
DICE 66.47 79.86 92.99 86.48 99.99
Focal–Tversky 67.90 80.30 98.59 87.32 99.98
Component ablations.
We ablated PLUM’s span extractor (SE) and feedback loop (F). Both contribute independently and jointly yield the best results.
Feedback loop robustness.
While conditioning on past inputs may theoretically introduce error accumulation, manual inspection showed that even when early masks were incorrect, PLUM’s feedback loop still produced correct later masks by relying on text embeddings. The loop also reduced duplicate or overlapping predictions. For example, the model without feedback repeated the same region for a lizard’s foot and tail, while PLUM with feedback predicted them distinctly.
Computational cost of feedback loop.
Timing experiments show that mask prediction is dominated by the LLM’s forward pass: averaged over 1000 examples, the LLM forward pass took 2.47 seconds, decoding without a feedback loop took .0097 seconds, and decoding with a feedback loop took .0162 seconds, a .2% increase compared to the language model’s time for its forward pass.
Loss weight sensitivity.
We varied the loss weights to verify stability with respect to $\lambda_{\mathrm{KL}}$ and $\lambda_{\mathrm{seg}}$. Performance changed modestly, confirming robustness.
Table 14: Loss weight sensitivity. PLUM is stable; slightly higher $\lambda_{\mathrm{KL}}$ improves performance.
$\lambda_{\mathrm{KL}}$ $\lambda_{\mathrm{seg}}$ $\lambda_{\mathrm{BCE}}$ LR gIoU cIoU
2.0 2.0 0.1 3e-4 0.385 0.331
4.0 2.0 0.1 3e-4 0.489 0.506
5.0 2.0 0.1 3e-4 0.517 0.494
6.0 2.0 0.1 3e-4 0.527 0.501
8.0 2.0 0.1 3e-4 0.573 0.546
Overall, these ablations confirm that PLUM’s architectural innovations—bidirectional span extraction, mask feedback conditioning, and Focal–Tversky optimization—jointly account for its strong fine-grained segmentation and reasoning capabilities with minimal computational overhead.
C.5 Partonomy-Core Samples
Fig.˜11 provides an example of each question type for the Explanatory Part Segmentation task from Partonomy-Core.
(a) An Identification question.
(a) An Intersection question.
(a) An Difference question.
(a) A Whole-to-Part question.
(a) A Part-to-Whole question.
Figure 11: Examples of the different question types from the Partonomy-Core dataset.